a sphere of mass two kg is resting inside a cube.The cube is moving  with velocity of 5t along x axis+2talong y axis in m/s.here t is in seconds.all surface are smooth.the sphere is at rest with respect to the cube.hat is totalforce exerted by the sphere on the cube

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
a(x) = dv(x)/dt = 5
a(y) = d v(y)/dt = 2
Now
Fx = m* a(x) = 2* 5 = 10
Fy = m* [a(y)+g] = 2*(2+10) = 24
F = \sqrt(F²x + F²y)
F = \sqrt100+576 = \sqrt676
F = 26 N
Regards
Arun (askIITians forum expert)